OpenStudy (anonymous):

How many solutions does x − 13x + 40 = have?

OpenStudy (anonymous):

1

OpenStudy (amistre64):

infinite

OpenStudy (anonymous):

x - 13x = 40 => -12x = 40 => x = ( 40 / -12 ) => x = ( 10 / -3 )

OpenStudy (anonymous):

you write down the rest amistre ;)

OpenStudy (amistre64):

where did you imagine this from: x - 13x = 40 ?

OpenStudy (amistre64):

x − 13x + 40 = any real number that you can think of, and then some prolly

OpenStudy (anonymous):

ah yeah, misread the =, woops

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

i would have answered 0 though

OpenStudy (amistre64):

we could compromise and say, 1/0 lol

OpenStudy (anonymous):

I concur

OpenStudy (ybarrap):

Question needs repositioning